Nitrogen and hydrogen combine at high temperature in the presence of a catalyst to produce ammonia assume .240 mol of N2 and .746 mol of H2 are present initially. N2(g)+3H2(g)—>2NH3(g)
1
Expert's answer
2018-02-12T07:29:52-0500
Solution
The nitrogen would react completely with
0.240 mol N2 (3 mol H2/1 mol N2) = 0.720 mol H2,
producing ammonia in amount
0.240 mol N2 (2 mol NH3/1 mol N2) = 0.480 mol NH3,
leaving unreacted H2 in amount
0.746 mol H2 - 0.720 mol H2 = 0.026 mol H2.
Answer: 0.480 mol of NH3, produced, 0.026 mol of H2 remains.